# The twisted coarse Baum–Connes conjecture and relative hyperbolic groups

### v1

@@ 3 Filtered $C^{*}$-algebra and controlled $K$-theory
In this section, we recall the concept of filtered $C^{*}$-algebra and controlled $K$-theory from [GWY24]. Using these techniques, we are able to obtain a controlled Mayer–Vietoris argument for the twisted coarse Baum–Connes conjecture.
@@
A filtration on a $C^{*}$-algebra $A$ is a collection of self-adjoint subspaces $(A_{r})_{r\geq 0}$ of $A$ indexed by the non-negative real numbers that satisfies the following properties:
@@
In the case of Roe algebras, the filtartion is defined by propagation. We apply this framework to the obstruction algebras $C^{*}_{L,0}(M;\Gamma(X,\mathcal{A}))$ defined in Definition 2.18. For $r>0$, we define
@@
Let $A$ be a non-unital $C^{*}$-algebra with a filtration $(A_{r})_{r}$. Then for any $x\in K_{*}(A)$, there exists $r>0$ and an element $y\in K^{r,\frac{1}{8}}(A)$ such that $C(y)=x$. Moreover if $x,y\in K^{s,\frac{1}{8}}(A)$ for some $s>0$, then there exists $r>0$ such that $x=y$ in $K^{r,\frac{1}{8}}(A)$.
In the following, we shall recall some techniques for the computation of quantitative $K$-theory formulated in [GWY24]. Those techniques are crucial to reduce to computation of the $K$-theory of Roe algebra of relative hyperbolic group to that of peripheral subgroups.
@@
The collection $(I^{\omega},J^{\omega};A^{\omega})_{\omega\in\Omega}$ of pairs of ideals and $C^{*}$-algebras containing them is uniformly excisive if for any $r_{0},m_{0}\geq 0$ and $\epsilon>0$, there are $r\geq r_{0}$, $m\geq 0$, and $\delta>0$ such that:
@@
    for any $\omega\in\Omega$ and any $a\in(A^{\omega}\otimes\mathcal{K})_{r_{0}}$ of norm at most $m_{0}$, there exist elements $b\in(I^{\omega}\otimes\mathcal{K})_{r}$ and $c\in(J^{\omega}\otimes\mathcal{K})_{r}$ of norm at most $m$ such that $\|a-(b+c)\|<\epsilon$;
@@
    there exists $b\in I^{\omega}_{r}\otimes\mathcal{K}\cap J^{\omega}_{r}\otimes\mathcal{K}$ such that $\|a-b\|<\epsilon$.
Theorem 3.10 (Proposition 7.7 [GWY24]).
@@ 4 Relative hyperbolic groups, decomposition complexity and the strategy of the proof
In this section, we recall some properties of relative hyperbolic groups from [Osi05] and outline the strategy of the proof of Theorem 1.3. In this section, we assume that $G$ is a finitely generated group which is hyperbolic relative to a finite family of subgroups $\{H_{1},\cdots,H_{n}\}$. We fix a symmetric generating set $S$ of $G$ and denote by $d_{S}$ and $d_{s\cup\bigcup H_{i}}$ the Cayley distance on $G$ with respect to the generating set $S$ and $S\cup\bigcup H_{i}$, respectively. We denote the ball of radius $n$ with respect to $d_{s\cup\bigcup H_{i}}$ centered at $e$ by $B(n)$.
@@
In this setting, the metric space $(G,d_{s\cup\bigcup H_{i}})$ is hyperbolic.
The next lemma plays the crucial role in the proof of the twisted Baum–Connes conjecture of $B(n)$. By definition of $B(n)$, we have
@@
Here we recall the notion of decomposition of metric spaces from [GTY12] and [GTY13].
@@
where each $X_{i,j}\in\mathcal{Y}$ and the distance between distinct $X_{i,j}$ and $X_{i,j^{\prime}}$ is grater than $r$ for $i=0,1$. A metric family $\mathcal{X}$ is $r$-decomposable over $\mathcal{Y}$ if each $X\in\mathcal{X}$ admits an $r$-decomposition over $\mathcal{Y}$.
@@
Assume each peripheral subgroup $H_{i}$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras. If $\mathcal{X}$ is a metric family consists of subspaces of $(G,d_{S})$ whose diameter with respect to the distance $d_{s\cup\bigcup H_{i}}$ is uniformly bounded, then $\mathcal{X}\in\mathfrak{D}$.
@@
We prove Lemma 4.7 in Section 5.3 and Lemma 4.8 in Section 5.2.
@@
In this section, we prove Lemma 4.7 and Lemma 4.8 to complete the proof of Theorem 1.3. The lemmas are consequences of controlled Mayer-Vietoris sequences, which first appeared in [Yu98]. The proof is divided into three parts. In the first part, we show that a decomposition of metric spaces induces a decomposition of $C^{*}$-algebras so that we can use the controlled Mayer-Vietoris sequence formulated in [GWY24]. Second, using the cutting-and-pasting argument, we reduce the problem to the twisted coarse Baum–Connes conjecture for $B(n)$ for each $n$. Finally, we show that any metric family consisting of subspaces of $B(n)$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras again using the cutting-and-pasting argument.
@@ 5.1 Decomposition
Let $(X,d)$ be a metric space and assume that $Y_{0},Y_{1}\subset X$ satisfies $X=Y_{0}\cup Y_{1}$. For each $r>0$ and $i=0,1$, we consider the equivalence relation on $Y_{i}$ generated by
$x\sim y\Leftrightarrow d(x,y)<3r$
and decompose $Y_{i}=\bigsqcup_{j}Y_{i}^{(j)}$ into $3r$-separated equivalence classes. On each disjoint union $\bigsqcup_{j}N_{r}(Y_{i}^{(j)})$, we consider the distance
$d_{i}(x,y)=\left\{\begin{array}[]{cc}d(x,y)&x\in N_{r}(Y_{i}^{(j)}),~y\in N_{r}(Y_{i}^{(j)})\text{ for some }j\\ \infty&\text{otherwise.}\end{array}\right$.
We denote by $Y_{i}^{+r}$ the space $\bigsqcup_{j}N_{r}(Y_{i}^{(j)})$ equipped with the distance $d_{i}$.
@@
Fix $s_{0}\geq 1$ and for each decomposition $\omega=(Y_{0},Y_{1})$ of $X$, define filtered subspaces by
$B_{r}^{\omega}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r});\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s_{0}}(Y_{1}^{+r});\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}$
$I_{r}^{\omega}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r});\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}$
$J_{r}^{\omega}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{1}^{+r});\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}$
and define a $C^{*}$-subalgebra $B^{\omega}:=\overline{\bigcup_{r\geq 0}B_{r}^{\omega}}$ of $C^{*}_{L,0}(P_{s_{0}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ and two ideals $I^{\omega}:=\overline{\bigcup_{r\geq 0}I_{r}^{\omega}(X)}$ and $J^{\omega}:=\overline{\bigcup_{r\geq 0}J_{r}^{\omega}(X)}$ of $B^{\omega}$.
@@
The family of triples $(I^{\omega},J^{\omega};B^{\omega})_{\omega}$ is uniformly excisive (Definition 3.9), where $\omega$ runs all decompositions $\omega=(Y_{0},Y_{1})$ of $X$.
@@
We prove the first condition in Definition 3.9. Decompose $P_{s_{0}}(Y_{0}^{+r})\cup P_{s_{0}}(Y_{1}^{+r})$ into $Z_{0}:=P_{s_{0}}(Y_{0}^{+r})\setminus P_{s_{0}}(Y_{1}^{+r})$, $Z_{1}:=P_{s_{0}}(Y_{1}^{+r})\setminus P_{s_{0}}(Y_{0}^{+r})$ and $Z_{0,1}:=P_{s_{0}}(Y_{0}^{+r})\cap P_{s_{0}}(Y_{1}^{+r})$. For any $A\in B_{r}^{\omega}$, define
$A_{0}$ $:=\chi_{Z_{0}}A\chi_{Z_{0}}+\chi_{Z_{0}}A\chi_{Z_{0,1}}+\chi_{Z_{0,1}}A\chi_{Z_{0}}+\chi_{Z_{0,1}}A\chi_{Z_{0,1}}$
@@
Then $A=A_{0}+A_{1}$, $A_{0}$ and $A_{1}$ have propagation less than $r$, $A_{0}$ is supported on $P_{s_{0}}(Y_{0}^{+r})\times P_{s_{0}}(Y_{0}^{+r})$ and $A_{1}$ is supported on $P_{s_{0}}(Y_{1}^{+r})\times P_{s_{0}}(Y_{1}^{+r})$. Moreover we have
@@
For any partial translation $v$ on $Y_{0}^{+r}$, the functions $(A_{0})^{v}$ is in $\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K})$ since it is a restriction of $A^{v}$ by the definition of the coarse algebra (9). Therefore $A_{0}\in I_{r_{0}}^{\omega}$. Similarly, we have $A_{1}\in J_{r_{0}}^{\omega}$.
Next, we prove the second condition. For $r>0$, we define a partition of unity $\{\mu_{0}^{(r)},\mu_{1}^{(r)}\}$ of $P_{s_{0}}(Y_{0}^{+r})\cup P_{s_{0}}(Y_{1}^{+r})$ by
$\mu_{i}^{(r)}(x):=\frac{d\left(x,{P_{s_{0}}(Y_{i}^{+r})}^{C}\right)}{d\left(x,{P_{s_{0}}(Y_{0}^{+r})}^{C}\right)+d\left(x,{P_{s_{0}}(Y_{1}^{+r})}^{C}\right)}.$
Since
$d\left(x,{P_{s_{0}}(Y_{1}^{+r})}^{C}\right)+d\left(x,{P_{s_{0}}(Y_{2}^{+r})}^{C}\right)\geq r,$
and
$d(x,y)\leq r_{0}\Rightarrow\left|d\left(x,{P_{s_{0}}(Y_{i}^{+r})}^{C}\right)-d\left(y,{P_{s_{0}}(Y_{i}^{+r})}^{C}\right)\right|\leq r_{0},$
it is easy to see that
$d(x,y)\leq r_{0}\Rightarrow\left|\mu_{i}^{(r)}(x)-\mu_{i}^{(r)}(y)\right|\leq\frac{r_{0}}{r}.$
Let $a\in I^{\omega}\otimes\mathcal{K}\cap J^{\omega}\otimes\mathcal{K}$ such that $d(a,a_{1})<\delta$ and $d(a,a_{2})<\delta$ for some $a_{1}\in(I^{\omega}\otimes\mathcal{K})_{r_{0}}$ and $a_{2}\in(J^{\omega}\otimes\mathcal{K})_{r_{0}}$. Then,
$\left\|[\mu_{i}^{(r)},a_{j}]_{x,y}\right\|=\left\|(a_{j})_{x,y}\left(\mu_{i}^{(r)}(x)-\mu_{i}^{(r)}(y)\right)\right\|\leq\|a_{j}\|\cdot\left\|\mu_{i}^{(r)}(x)-\mu_{i}^{(r)}(y)\right\|.$
Fix any $\varepsilon>0$. By [Roe03, Lemma 4.27], there exists a constant $C(r_{0})$ such that
$\left\|[\mu_{i}^{(r)},a_{j}]\right\|\leq C(r_{0})\frac{r_{0}}{r}\|a_{j}\|<\frac{\varepsilon}{4}$
by taking $r>\frac{4r_{0}C(r_{0})}{\varepsilon}\|a_{j}\|$. Now define
$b:=\sqrt{\mu_{2}}a_{1}\sqrt{\mu_{2}}+\sqrt{\mu_{1}}a_{2}\sqrt{\mu_{1}}.$
Then $b$ is supported on $\left(P_{s_{0}}(Y_{0}^{+r})\times P_{s_{0}}(Y_{0}^{+r})\right)\cap\left(P_{s_{0}}(Y_{1}^{+r})\times P_{s_{0}}(Y_{1}^{r})\right)$. Therefore, by the third condition of Definition 2.1, we have
$b\in I^{\omega}_{r}\otimes\mathcal{K}\cap J^{\omega}_{r}\otimes\mathcal{K}.$
In addition, if $\delta<\frac{\varepsilon}{4}$, we have
$\|a-b\|$
$\leq$ $\|a-\left(\sqrt{\mu_{2}}a\sqrt{\mu_{2}}+\sqrt{\mu_{1}}a\sqrt{\mu_{1}}\right)\|+\|(\sqrt{\mu_{2}}a\sqrt{\mu_{2}}+\sqrt{\mu_{1}}a\sqrt{\mu_{1}})-(\sqrt{\mu_{2}}a_{1}\sqrt{\mu_{2}}+\sqrt{\mu_{1}}a_{2}\sqrt{\mu_{1}})\|$
$\leq$ $\frac{\varepsilon}{2}+2\delta<\varepsilon$

### v2

@@ 3 Filtered $C^{*}$-algebra and controlled $K$-theory
In this section, we recall the concepts of filtered $C^{*}$-algebras and controlled $K$-theory from [GWY24]. Using these techniques, we obtain a controlled Mayer–Vietoris argument for the twisted coarse Baum–Connes conjecture.
@@
A filtration on a $C^{*}$-algebra $A$ is a collection of self-adjoint subspaces $(A_{r})_{r\geq 0}$ of $A$ indexed by the non-negative real numbers with the following properties:
@@
In the case of Roe algebras, the filtration is defined by propagations. We apply this framework to the obstruction algebras $C^{*}_{L,0}(M;\Gamma(X,\mathcal{A}))$ defined in Definition 2.18. For $r>0$, we define
@@
Let $A$ be a non-unital $C^{*}$-algebra with a filtration $(A_{r})_{r}$. Then for any $x\in K_{*}(A)$, there exists $r>0$ and an element $y\in K^{r,\frac{1}{8}}(A)$ such that $c(y)=x$. Moreover if $x,y\in K_{*}^{s,\frac{1}{8}}(A)$ for some $s>0$ satisfies $c(x)=c(y)$, then there exists $r>0$ such that $x=y$ in $K_{*}^{r,\frac{1}{8}}(A)$.
In the following, we shall recall some techniques for the computation of quantitative $K$-theory formulated in [GWY24]. Those techniques are crucial to reduce to computation of the $K$-theory of twisted Roe algebras of relative hyperbolic group to that of peripheral subgroups.
@@
Now, we recall the following techniques to compute the quantitative $K$-theory.
@@
The collection $(I^{\omega},J^{\omega};A^{\omega})_{\omega\in\Omega}$ of pairs of ideals and $C^{*}$-algebras containing them is uniformly excisive if for any $r_{0},m_{0}\geq 0$ and $\varepsilon>0$, there are $r\geq r_{0}$, $m\geq 0$, and $\delta>0$ such that:
@@
    for any $\omega\in\Omega$ and any $a\in(A^{\omega}\otimes\mathcal{K})_{r_{0}}$ of norm at most $m_{0}$, there exist elements $b\in(I^{\omega}\otimes\mathcal{K})_{r}$ and $c\in(J^{\omega}\otimes\mathcal{K})_{r}$ of norm at most $m$ such that $\|a-(b+c)\|<\varepsilon$;
@@
    there exists $b\in I^{\omega}_{r}\otimes\mathcal{K}\cap J^{\omega}_{r}\otimes\mathcal{K}$ such that $\|a-b\|<\varepsilon$.
Theorem 3.10 (Proposition 7.7, [GWY24]).
@@ 4 Relative hyperbolic groups, decomposition complexity and the strategy of the proof
In this section, we recall some properties of relative hyperbolic groups from [OSI05] and outline the strategy of the proof of Theorem 1.3. In this section, we assume that $G$ is a finitely generated group which is hyperbolic relative to a finite family of subgroups $\{H_{1},\cdots,H_{n}\}$.
We fix a symmetric generating set $S$ of $G$ and denote by $d_{S}$ and $d_{S\cup\bigcup H_{i}}$ the Cayley distance on $G$ with respect to the generating set $S$ and $S\cup\bigcup H_{i}$, respectively. We denote by $B(n)$ the ball of radius $n$ under the metric $d_{S\cup\bigcup H_{i}}$ on $G$.
@@
In the above setting, the metric space $(G,d_{S\cup\bigcup H_{i}})$ is hyperbolic.
By definition of $B(n)$, we have
@@
The following lemma plays a crucial role in the proof of the twisted Baum–Connes conjecture of $B(n)$.
@@
Recall that a family of subspaces $\{X_{i}\}_{i}$ of a metric space $(X,d)$ is said to be $L$-separated if $d(X_{i},X_{j})\geq L$ for $i\neq j$.
Now, we recall the notion of decomposition of metric spaces from [GTY12] and [GTY13].
@@
where each $X_{i,j}\in\mathcal{Y}$ and for $i=0,1$, the metric family $\{X_{i,j}\}_{j}$ is $r$-separated. A metric family $\mathcal{X}$ is $r$-decomposable over $\mathcal{Y}$ if each $X\in\mathcal{X}$ admits an $r$-decomposition over $\mathcal{Y}$.
@@
Assume that each peripheral subgroup $H_{i}$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras. If $\mathcal{X}$ is a metric family consists of subspaces of $(G,d_{S})$ whose diameter with respect to the distance $d_{s\cup\bigcup H_{i}}$ is uniformly bounded, then $\mathcal{X}\in\mathfrak{D}$.
@@
We will prove Lemma 4.7 in Section 5.3 and Lemma 4.8 in Section 5.2.
Combining Lemmas 4.7 and 4.8, we obtain our main result.
Theorem 4.9.
@@
In this section, we prove Lemma 4.7 and Lemma 4.8 to complete the proof of Theorem 1.3. The lemmas are consequences of controlled Mayer-Vietoris sequences, which first appeared in [YU98]. The proof is divided into three parts. In the first part, we show that a decomposition of metric spaces induces a decomposition of obstruction algebras so that we can use the controlled Mayer-Vietoris sequence formulated in [GWY24]. Second, using the cutting-and-pasting argument, we reduce the problem to the twisted coarse Baum–Connes conjecture for $B(n)$ for each $n$. Finally, we show that any metric family consisting of subspaces of $B(n)$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras again using the cutting-and-pasting argument.
@@ 5.1 Decomposition
Fix $s_{0}\in\mathbb{N}$. Let $(X,d)$ be a uniformly locally finite metric space and assume that $\omega=(Y_{0},Y_{1})$ be a decomposition of $X$. (i.e. $Y_{0}$ and $Y_{1}$ are subspaces of $X$ such that $X=Y_{0}\cup Y_{1}$.) Let $R_{r}$ be an increasing function of $r$ such that $R_{r}>(s_{0}+1)r$. We decompose $Y_{i}=\bigsqcup Y_{i}^{j}$ into $3R_{r}$-separated subsets. Then for $i=0,1$ we have
$N_{r}(P_{s_{0}}(Y_{i}))=\bigsqcup_{j}N_{r}(P_{s_{0}}(Y_{i}^{j})),$
where $N_{r}$ is the $r$-neighborhood in $P_{s_{0}}(X)$ and the disjoint union is at least $R_{r}$-separated. We define the new distance $d^{R,r}_{i}$ on $N_{r}(P_{s_{0}}(Y_{i}))=\bigsqcup_{j}N_{r}(P_{s_{0}}(Y_{i}^{j}))$ by
$d_{i}^{R,r}(x,y)=\left\{\begin{array}[]{cc}d(x,y)&x\in N_{r}(P_{s_{0}}(Y_{i}^{j})),~y\in N_{r}(P_{s_{0}}(Y_{i}^{j}))\text{ for some }j\\ \infty&\text{otherwise.}\end{array}\right$.
We denote by $P_{s_{0}}(Y_{i})^{+r}$ the space $N_{r}(P_{s_{0}}(Y_{i}))$ equipped with the distance $d_{i}^{R,r}$.
@@
We first define filtered subspaces of $C^{*}_{L,0}(P_{s_{0}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ by
$B_{r}^{X,R,\omega}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0})^{+r};\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s_{0}}(Y_{1})^{+r};\Gamma(Y_{1},\mathcal{A}\otimes\mathcal{K}))_{r},$
$I_{r}^{X,R,\omega}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0})^{+r};\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K}))_{r},$
$J_{r}^{X,R,\omega}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{1})^{+r};\Gamma(Y_{1},\mathcal{A}\otimes\mathcal{K}))_{r}.$
Then we define $C^{*}$-subalgebras of $C^{*}_{L,0}(P_{s_{0}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$:
$B^{X,R,\omega}:=\overline{\bigcup_{r\geq 0}B_{r}^{\omega}},~I^{X,R,\omega}:=\overline{\bigcup_{r\geq 0}I_{r}^{\omega}(X)},~\mbox{and}~I^{X,R,\omega}:=\overline{\bigcup_{r\geq 0}I_{r}^{\omega}(X)}.$
Now, we show that $I^{X,R,\omega}$ and $J^{X,R,\omega}$ are filtered ideals of $B^{X,R,\omega}$ (see Definition 3.8). It suffices to show that for any $S\in C^{*}(P_{s_{0}}(Y_{1})^{+r_{1}};\Gamma(Y_{1},\mathcal{A}\otimes\mathcal{K}))_{r_{1}}$ and $T\in C^{*}(P_{s_{0}}(Y_{0})^{+r_{2}};\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K}))_{r_{2}}$, the product $ST$ is an element in $C^{*}(P_{s_{0}}(Y_{0})^{+r_{1}+r_{2}};\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K}))_{r_{1}+r_{2}}$. We only need to verify the condition (1) of Definition 2.7. Let us fix a Borel cover $\{B_{x}\}_{x\in Y_{0}}$ of $P_{s_{0}}(Y_{0})^{+r_{1}+r_{2}}$ satisfying the four conditions in Lemma 2.4. For any partial translation $v$ on $Y_{0}$, we show that the function
$(ST)^{v}:Y_{0}\to\mathcal{A}\otimes\mathcal{K},~x\mapsto\chi_{B_{v(x)}}(ST)\chi_{B_{x}}=\sum_{y}\left(\chi_{B_{v(x)}}S\chi_{B_{y}}\right)\left(\chi_{B_{y}}T\chi_{B_{x}}\right)$
is in $\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K})$. Both $S$ and $T$ can be decomposed into finite sums of the operators of the form $f\cdot v$, where $f$ is an $\ell^{\infty}$ function and $v$ is a partial translation. By further decomposing partial translations, we may assume that for each $x\in Y_{0}$, there is at most one $y$ such that $\chi_{B_{v(x)}}S\chi_{B_{y}}\neq 0$ and $\chi_{B_{y}}T\chi_{B_{x}}\neq 0$. We denote such $y$ by $\alpha(x)$ and denote the assignment $v(x)\to y$ by $\beta$, and then $\alpha$ and $\beta$ are partial translations on $Y_{0}$. In this case, we have that
$(ST)^{v}(x)=\left(\chi_{B_{v(x)}}S\chi_{B_{\beta v^{-1}(x)}}\right)\left(\chi_{B_{\alpha(x)}}T\chi_{B_{x}}\right).$
By the definition of coarse algebras for subspaces (9), there are $\varphi,\psi\in\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K})$ such that $\varphi(x)=\chi_{B_{v(x)}}S\chi_{B_{\beta v^{-1}(x)}}$ and $\psi(x)=\chi_{B_{\alpha(x)}}T\chi_{B_{x}}$ for all $x\in Y_{0}$. Therefore the restriction of $\varphi\psi$ to $Y_{0}$ is also an element in $\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K})$.
@@
The family of triples $(I^{X,R,\omega},J^{X,R,\omega};B^{X,R,\omega})_{X,R,\omega}$ is uniformly excisive (Definition 3.9), where $X$ runs all metric space with bounded geometry, $\omega$ runs over all decompositions $\omega=(Y_{0},Y_{1})$ of $X$ and $R$ runs over all increasing functions of $r$ satisfying $R_{r}>(3s_{0}+1)r$.
@@
For given $m_{0},\varepsilon,r_{0}$ in Definition 3.9, we prove that $m=4m_{0}$, $\delta=\frac{1}{2}\varepsilon$ and $r=r_{0}+1$ work.
First, we prove the first condition of Definition 3.9. Note that
$P_{s_{0}}(X)\subset N_{1}(P_{s_{0}}(Y_{0}))\cup N_{1}(P_{s_{0}}(Y_{1})).$
We define
$Z_{0}$ $:=N_{r_{0}}(P_{s_{0}}(Y_{0}))\setminus N_{r_{0}}(P_{s_{0}}(Y_{1})),$
$Z_{1}$ $:=N_{r_{0}}(P_{s_{0}}(Y_{1}))\setminus N_{r_{0}}(P_{s_{0}}(Y_{0})),$
$Z_{0,1}$ $:=N_{r_{0}}(P_{s_{0}}(Y_{0}))\cap N_{r_{0}}(P_{s_{0}}(Y_{1})).$
Note that $d(Z_{0},Z_{1})>r_{0}$. For any $A\in B_{r_{0}}^{X,R,\omega}$, we define
$A_{0}$ $:=\chi_{Z_{0}}A\chi_{Z_{0}}+\chi_{Z_{0}}A\chi_{Z_{0,1}}+\chi_{Z_{0,1}}A\chi_{Z_{0}}+\chi_{Z_{0,1}}A\chi_{Z_{0,1}},$
@@
Since $A$ has propagation less than $r_{0}$, we have $A=A_{0}+A_{1}$. In addition, $A_{0}$ and $A_{1}$ have propagation less than $r_{0}$, $A_{0}$ is supported on $N_{r_{0}}(P_{s_{0}}(Y_{0}))\times N_{r_{0}}(P_{s_{0}}(Y_{0}))$ and $A_{1}$ is supported on $N_{r_{0}}(P_{s_{0}}(Y_{1}))\times N_{r_{0}}(P_{s_{0}}(Y_{1}))$. Moreover we have
@@
For any partial translation $v$ on $Y_{0}$, the functions $(A_{0})^{v}$ is in $\Gamma(Y_{0},\mathcal{A}\otimes\mathcal{K})$ by the definition of the coarse algebra (9) since it is a restriction of $A^{v}$. Therefore $A_{0}\in I_{r_{0}}^{\omega}$. Similarly, we have $A_{1}\in J_{r_{0}}^{\omega}$.
Next, we prove the second condition. Let $\{\mu_{0},\mu_{1}\}$ be a partition of unity subordinate to the cover $N_{1}(P_{s_{0}}(Y_{0}))\cup N_{1}(P_{s_{0}}(Y_{1}))$. Let $A\in I^{X,R,\omega}\otimes\mathcal{K}\cap J^{X,R,\omega}\otimes\mathcal{K}$ be such that $\|A-A_{0}\|\leq\delta$ and $\|A-A_{1}\|\leq\delta$ for some $A_{0}\in\left(I^{X,R,\omega}\right)_{r_{0}}$ and $A_{1}\in\left(J^{X,R,\omega}\right)_{r_{0}}$. Define
$B:=A_{0}\mu_{1}+A_{1}\mu_{0}.$
Then since the multiplication by $\mu_{1}$ does not increase the support,
$\operatorname{supp}{(A_{0}\mu_{1})}\subset\operatorname{supp}(A_{0})\subset N_{r_{0}}(P_{s_{0}}(Y_{0}))\times N_{r_{0}}(P_{s_{0}}(Y_{0})).$
Since $\mu_{1}$ is supported on $N_{1}(P_{s_{0}}(Y_{1}))$, and $A_{0}$ has propagation less than $r_{0}$, we obtain that
$\operatorname{supp}{(A_{0}\mu_{1})}\subset N_{r_{0}+1}(P_{s_{0}}(Y_{1}))\times N_{1}(P_{s_{0}}(Y_{1}))\subset N_{r_{0}+1}(P_{s_{0}}(Y_{1}))\times N_{r_{0}+1}(P_{s_{0}}(Y_{1}))$
By arguing similarly for the support of $A_{1}\mu_{0}$, we can show that
$\operatorname{supp}(B)\subset\left(N_{r_{0}+1}(P_{s_{0}}(Y_{0}))\cap N_{r_{0}+1}(P_{s_{0}}(Y_{1}))\right)\times\left(N_{r_{0}+1}(P_{s_{0}}(Y_{0}))\cap N_{r_{0}+1}(P_{s_{0}}(Y_{1}))\right).$
Also, we have
$\|A-B\|=\|(A-A_{0})\mu_{1}+(A-A_{1})\mu_{0}\|\leq 2\varepsilon=\delta.$